for the data in the table tell whether y varies directly with x if it does write an equation for the direct variation
x y
0 0
1 -2.5
2 -5
3 -7.5

Found 3 solutions by mananth, ewatrrr, stanbon:
Answer by mananth(16946)

x y
0 0
1 -2.5
2 -5
3 -7.5
..
y = 2.5x is the equation

Answer by ewatrrr(24785)


Hi,
*Note: when x = 0, then y = 0
Letting k represent the constant to see if y varies directly with x:
y = k*x
using Pt (1,(-5/2)) to solve for k
(-5/2) = k*1
(-5/2) = k
checking k against the remaining data
-5 = (-5/2)*2
-7.5 = -7 1/2 = (-15/2)= (-5/2)*3
y = (-5/2)x is the equation for this direct variation

Answer by stanbon(75887)

for the data in the table tell whether y varies directly with x if it does write an equation for the direct variation
x y
0 0
1 -2.5
2 -5
3 -7.5
-------------
(0,0) tells you the y-intercept is 0
---
Since y decreases by 2.5 for each x-increase of one
the slope is -2.5
----
Equation:
y = -2.5x
==================
The variation is direct.
All linear function have direct variation.
------------------
Cheers,
Stan H.